OpenStudy (anonymous):

Evaluate the expression below when x = -2. x2 − 8 x + 6 -3 -1 1 2 3 2

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

substitute x for -2

Nnesha (nnesha):

\[\huge\rm \color{Red}{x}^2-8\] replace x with -2 and then solve!

OpenStudy (anonymous):

ok think you so much

Nnesha (nnesha):

wait i don't get it there are two expression x^2-8 and x+6 right ?

OpenStudy (anonymous):

ya

Nnesha (nnesha):

can you take a screenshot of ur question plz ?

OpenStudy (anonymous):

how do you do that

Nnesha (nnesha):

press \(\huge\color{green}{\rm PrntSCr}\) and then paste that into *paint * save it click on attach file button to post

OpenStudy (anonymous):

is that it

Nnesha (nnesha):

ohh division you should use */* to show something is dividing \[\frac{ x^2 -8}{ x+6 }\] now substitute x for -2

Nnesha (nnesha):

let me know what you get :)

OpenStudy (anonymous):

I got -2/4

Nnesha (nnesha):

how did you get -2 at the numerator ?

Nnesha (nnesha):

remember it's x^2 so you should take square of -2

OpenStudy (anonymous):

what are you talking about idk this stuff

OpenStudy (anonymous):

@triciaal

Nnesha (nnesha):

\[\huge\rm \frac{ \color{reD}{x}^2 -8}{ x+6 }\] substitute x for -2 put the parentheses

Nnesha (nnesha):

\[\huge\rm \frac{ (?)^2+8 }{ (?)+6 }\]do you see the question marks that's where you should put -2

OpenStudy (anonymous):

yes I see it but I need a calculator

Nnesha (nnesha):

you don't really need one.

Nnesha (nnesha):

\[\huge\rm \frac{ (-2)^2+8 }{ (-2)+6 }\] (-2)^2 is same as -2 times -2

OpenStudy (anonymous):

why do we do -2 times-2

Nnesha (nnesha):

(-2)^2 means -2 times -2

OpenStudy (anonymous):

ok I just got a calculator and I got 14

OpenStudy (anonymous):

@Elsa213

OpenStudy (anonymous):

@Hero

Nnesha (nnesha):

how did you get 14 ? what is -2 times -2 ?